# Isometric rigidity of the Wasserstein space over the plane with the maximum metric

### v1

@@ 3. $\bullet$
The second condition is the injectivity of the Radon transform with respect to the variable $\mu$ on the set $\mathcal{F}$. Suppose that we have such an $\mathcal{F}$. Then, applying Lemma 5 we get that for any $\mu\in\mathcal{F}$
@@
By the third condition we have $\Phi(\mu)\in\mathcal{F}$ and so we can apply the second condition for the two measures $\mu_{1}=\mu$ and $\mu_{2}=\Phi(\mu)$. This implies that $\Phi(\mu)=\mu$ for all $\mu\in\mathcal{F}$. Using the density of $\mathcal{F}$ in $\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})$ (the first condition) we get that $\Phi(\mu)=\mu$ for all $\mu\in W_{p}(\mathbb{R}^{2})$.
@@
$\mathcal{F}:=$$\Bigg{\{}\sum_{i=1}^{N}a_{i}\delta_{x_{i}}\colon N\geq 1,\sum_{i=1}^{N}a_{i}=1,$
@@
We would like to check the third property of $\mathcal{F}$ let us take an element $\mu\in\mathcal{F}.$ Note that by Lemma 5 we can write that
$\Phi({P_{L}}_{\#}(\mu))={P_{L}}_{\#}(\Phi(\mu))={P_{L}}_{\#}(\mu)$
for $L\in\{L_{+},L_{-}\}$. Let us denote by
${P_{L_{+}}}_{\#}(\Phi(\mu))={P_{L_{+}}}_{\#}(\mu)=\sum_{i=1}^{N}a_{i}\xi_{i},$
where $\xi_{i}\in L_{+}$ and
${P_{L_{-}}}_{\#}(\Phi(\mu))={P_{L_{-}}}_{\#}(\mu)=\sum_{i=1}^{N}a_{i}\eta_{i},$
@@
where $\eta_{i}\in L_{-}$. Now we can see that $\Phi(\mu)$ is a finitely supported measure with support contained in the intersection of the two pre-images:
$(P_{L_{+}})^{-1}(\{\xi_{1},\ldots,\xi_{N}\})\cap(P_{L_{-}})^{-1}(\{\eta_{1},\ldots,\eta_{N}\}).$
Making use of the condition that $a_{i}\neq a_{j}$ if $i\neq j$ we see that the only possibility for $\Phi(\mu)$ is to be the measure
$\Phi(\mu)=\sum_{i=1}^{N}a_{i}\delta_{x_{i}},$
where $x_{i}=(P_{L_{+}})^{-1}(\xi_{i})\cap(P_{L_{+}})^{-1}(\eta_{i})$, proving that $\mathcal{F}$ satisfies the third condition as well.
@@ 3.2. Diagonal rigidity of $\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})$ for $p>1$
To handle the case $\Phi_{t}$ for $t>0$, choose $\mu=\mu(0,1,0)=\frac{1}{2}\delta_{(-1,-1)}+\frac{1}{2}\delta_{(1,1)}$. Then $\Phi(\mu(0,1,0))=\mu(0,1,t)$. Now fix the Dirac measure $\delta_{(-1,0)}$ and calculate $d_{W_{2}}(\delta_{(-1,0)},\mu(0,1,0)$ and $d_{W_{2}}(\delta_{(-1,0)},\mu(0,1,t)$. Again, if $\Phi$ would act like $\Phi_{t}$ on $L_{+}$, we should get the same result, since $\Phi$ is an isometry. The calculation shows that $d_{W_{2}}\big{(}\delta_{(-1,0)},\mu(0,1,0)\big{)}=\sqrt{\frac{5}{2}}$ and
@@ 3.3. Diagonal rigidity of $\mathcal{W}_{1}(Q,d_{\mathrm{{m}}})$
In this subsection, we consider the case when $X=Q=[0,1]^{2}$ and $p=1$. Diagonal rigidity is achieved as a result of the following statements about measures that are supported on the sides, at the corners, and finally on the diagonals of $Q$. The first statement concerns measures supported on the opposite sides of the boundary of $Q$ and it is valid for all $p\geq 1$.
@@
Let $p\geq 1$ and $\Phi:\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})\to\mathcal{W}_{p}(Q,d_{\mathrm{{m}}}),d_{W_{p}})$ be an isometry. If $\mu,\nu\in\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})$ are two probability measures whose supports lie on opposite sides of the unit square $Q=[0,1]^{2}$, then their isometric images $\Phi(\mu)$ and $\Phi(\nu)$ have the same property.
@@
with equality if and only if the supports $\mathrm{supp}(\mu)$ and $\mathrm{supp}(\nu)$ are contained in two opposite sides of $Q$. To see this, note that inequality (29) follows immediately from the definition of the Wasserstein metric $d_{W_{p}}$ and the fact that $d_{\mathrm{{m}}}(x,y)\leq 1$ for all $x,y\in Q$.
Furthermore note, that if $\mathrm{supp}(\mu)$ and $\mathrm{supp}(\nu)$ are contained in two opposite sides of $Q$, then $d_{\mathrm{{m}}}(x,y)=1$ for all $x\in\mathrm{supp}(\mu)$ and $y\in\mathrm{supp}(\nu)$.
Let $\pi_{0}$ be an optimal coupling of $\mu$ and $\nu$. Since $\mathrm{supp}(\pi_{0})\subseteq\mathrm{supp}(\mu)\times\mathrm{supp}(\nu)$ we have that $d_{\mathrm{{m}}}(x,y)=1$ for any $(x,y)\in\mathrm{supp}(\pi_{0})$ and therefore
$d_{W_{p}}^{p}(\mu,\nu)=\int\limits_{Q}d_{\mathrm{{m}}}^{p}(x,y)~{}\mathrm{d}\pi_{0}(x,y)=1.$
To show the converse, let us assume that $d_{W_{p}}^{p}(\mu,\nu)=1$ and assume by contradiction that $\mathrm{supp}(\mu)$ and $\mathrm{supp}(\nu)$ do not lie on two opposite sides of $Q$. This implies that there is a point $x_{0}\in\mathrm{supp}(\mu)$ with the property that $d_{\mathrm{{m}}}(x_{0},y)<1$ for any $y\in\mathrm{supp}(\nu)$. We can even find a small radius $r>0$ and $\varepsilon>0$ such that
$d_{\mathrm{{m}}}(x,y)<1-\varepsilon\ \text{for all}\ x\in B(x_{0},r)\cap Q\ \text{and}\ y\in\mathrm{supp}(\nu).$
Since $x_{0}\in\mathrm{supp}(\mu)$ we have that $\delta:=\mu(B(x_{0},r))>0$. Let us consider the trivial coupling $\pi=\mu\otimes\nu$. Then we have
$\begin{split}&d_{W_{p}}^{p}(\mu,\nu)\leq\int\limits_{\mathrm{supp}(\mu)}\int\limits_{\mathrm{supp}(\nu)}d_{\mathrm{{m}}}^{p}(x,y)~{}\mathrm{d}\mu(x)~{}\mathrm{d}\nu(y)=\\ &=\int\limits_{\mathrm{supp}(\mu)\cap B(x_{0},r)}\int\limits_{\mathrm{supp}(\nu)}d_{\mathrm{{m}}}^{p}(x,y)~{}\mathrm{d}\mu(x)~{}\mathrm{d}\nu(y)+\int\limits_{\mathrm{supp}(\mu)\setminus B(x_{0},r)}\int\limits_{\mathrm{supp}(\nu)}d_{\mathrm{{m}}}^{p}(x,y)~{}\mathrm{d}\mu(x)~{}\mathrm{d}\nu(y)\\ &\leq\ (1-\varepsilon)^{p}\delta+1(1-\delta)<1.\end{split}$
This contradicts the fact $d_{W_{p}}^{p}(\mu,\nu)=1$. This proves the converse implication, namely that if $d_{W_{p}}^{p}(\mu,\nu)=1$ then the supports of $\mu$ and $\nu$ must be contained in opposite sides of $Q$.
The statement of the lemma is now an immediate consequence of this claim. Indeed, assume that $\mathrm{supp}(\mu)$ and $\mathrm{supp}(\nu)$ are contained in two opposite sides of $Q$. Then we have $d_{W_{p}}^{p}(\mu,\nu)=1$. Since $\Phi:\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})\to\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})$ is an isometry, we have $d_{W_{p}}^{p}(\Phi(\mu),\Phi(\nu))=1.$ But, then the supports of the two measures $\Phi(\mu)$ and $\Phi(\nu)$ must be contained in two opposite sides of $Q$. ∎
@@
$V=\{\delta_{x_{1}}=\delta_{(0,0)},\delta_{x_{2}}=\delta_{(1,0)},\delta_{x_{3}}=\delta_{(1,1)},\delta_{x_{4}}=\delta_{(0,1)}\}.$
@@
It is clear by Lemma 12 that all other measures $\Phi(\delta_{x_{i}})$ for $i=2,3,4$ must be supported on the opposite side of $S_{1}$, that is $S_{3}.$ But the mutual distance of any pair of these measures must be equal to $1$ which shows that any two of these three measures must be lying on opposite sides again, which is $S_{2}$ and $S_{4}$. But there are only two possibilities for measures with support in $S_{2}\cap S_{3}$ and $S_{4}\cap S_{3}$, namely the two Dirac masses on the vertices of $S_{3}$ which gives a contradiction.

### v3

@@ 3. $\bullet$
The second condition is the injectivity of the Radon transform on the set $\mathcal{F}$. Suppose that we have such an $\mathcal{F}$. Then, applying Lemma 6 we get that for any $\mu\in\mathcal{F}$
@@
By the third condition we have $\Phi(\mu)\in\mathcal{F}$ and so we can apply the second condition for the two measures $\mu_{1}=\mu$ and $\mu_{2}=\Phi(\mu)$. This implies that $\Phi(\mu)=\mu$ for all $\mu\in\mathcal{F}$. Using the density of $\mathcal{F}$ in $\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})$ (the first condition) we get that $\Phi(\mu)=\mu$ for all $\mu\in\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})$.
@@
$\mathcal{F}:=$$\Bigg{\{}\sum_{i=1}^{N}a_{i}\delta_{x_{i}}\,\Bigg{|}N\geq 1,\sum_{i=1}^{N}a_{i}=1,$
@@
To verify the third property, i.e. that $\Phi(\mathcal{F})\subseteq\mathcal{F}$, let us take an element $\mu=\sum_{i=1}^{N}a_{i}\delta_{x_{i}}\in\mathcal{F}$. Recalling that $\Phi$ fixes all measures supported on the diagonals $L_{+}$ and $L_{-}$, we get by Lemma 6 that
${P_{L}}_{\#}(\mu)=\Phi({P_{L}}_{\#}(\mu))={P_{L}}_{\#}(\Phi(\mu))$
for $L\in\{L_{+},L_{-}\}$, and therefore we have
${P_{L_{+}}}_{\#}(\Phi(\mu))={P_{L_{+}}}_{\#}(\mu)=\sum_{i=1}^{N}a_{i}\delta_{P_{L_{+}}(x_{i})},$
where $P_{L_{+}}(x_{i})\in L_{+}$ and
${P_{L_{-}}}_{\#}(\Phi(\mu))={P_{L_{-}}}_{\#}(\mu)=\sum_{i=1}^{N}a_{i}\delta_{P_{L_{-}}(x_{i})},$
where $P_{L_{-}}(x_{i})\in L_{-}$. In conclusion we obtain that the Radon transform of $\mu$ and $\Phi(\mu)$ are equal
$\mathcal{R}:=\mathcal{R}(\mu)=\mathcal{R}(\Phi(\mu))=\left(\sum_{i=1}^{N}a_{i}\delta_{P_{L_{+}}(x_{i})},\sum_{i=1}^{N}a_{i}\delta_{P_{L_{-}}(x_{i})}\right).$
@@
Now observe that $\Phi(\mu)$ is a finitely supported measure with support contained in the intersection of the two pre-images:
$(P_{L_{+}})^{-1}(\{P_{L_{+}}(x_{1}),\ldots,P_{L_{+}}(x_{N})\})\cap(P_{L_{-}})^{-1}(\{P_{L_{-}}(x_{1}),\ldots,P_{L_{-}}(x_{N})\}).$
This intersection is an $N$-by-$N$ grid, and we refer to its points by $z_{i,j}$ $(1\leq i,j\leq N)$, where $z_{i,j}$ satisfies that $P_{L_{+}}(z_{i,j})=P_{L_{+}}(x_{i})$ and $P_{L_{-}}(z_{i,j})=P_{L_{-}}(x_{j})$. Hence $\Phi(\mu)$ can be written as
$\Phi(\mu)=\sum_{i=1}^{N}\sum_{j=1}^{N}a_{i,j}\delta_{z_{i,j}},$
where $a_{i,j}\geq 0$, $\sum_{i=1}^{N}\sum_{j=1}^{N}a_{i,j}=1$ and $\sum_{i=1}^{N}a_{i,j}=a_{j}$ and $\sum_{j=1}^{N}a_{i,j}=a_{i}$. Since the grid is finite, there is a positive minimal distance between its points
$c:=\min_{(i,j)\neq(i^{\prime},j^{\prime})}d_{m}(z_{i,j},z_{i^{\prime},j^{\prime}})>0.$
From here, we assume by contradiction that $\Phi(\mu)\notin\mathcal{F}$. Since $\Phi(\mu)\notin\mathcal{F}$, then there exist at least two points $z,z^{\prime}\in\mathrm{supp}(\Phi(\mu))$ such that their projection onto either $L_{+}$ or $L_{-}$ coincide.
Figure 1. Illustration of a finitely supported measure $\mu$ with a possible image $\Phi(\mu)$ and the grid determined by the pre-images of $P_{L_{+}}$ and $P_{L_{-}}$.
We briefly sketch how to obtain the desired contradiction and we give the details later:
First, by slightly perturbing the measure $\mu$ we will construct a measure $\mu^{\prime}$ such that
(3) $\arg\min\big{\{}d_{W_{p}}(\mu,\xi)\,\big{|}\,\xi\in\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}}),\mathcal{R}(\xi)=\mathcal{R}(\mu^{\prime})\big{\}}=\{\mu^{\prime}\}.$
Figure 2. Illustration of $\mu^{\prime}$, the measure that we obtain by a sufficiently small perturbation of $\mu$.
Next, using the existence of $z$ and $z^{\prime}$ as above; by small perturbations of $\Phi(\mu)$ we will construct two measures $\nu_{1}^{\prime}$ and $\nu_{2}^{\prime}$ such that $\mathcal{R}(\nu_{1}^{\prime})=\mathcal{R}(\nu_{2}^{\prime})=\mathcal{R}(\mu^{\prime})$ and
(4) $\arg\min\big{\{}d_{W_{p}}(\Phi(\mu),\xi)\,\big{|}\,\xi\in\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}}),\mathcal{R}(\xi)=\mathcal{R}(\mu^{\prime})\big{\}}\supseteq\{\nu_{1}^{\prime},\nu_{2}^{\prime}\}.$
Figure 3. Illustration for $\nu_{1}^{\prime}$ and $\nu_{2}^{\prime}$ the two measures that we obtained by sufficiently small perturbations of $\Phi(\mu)$.
Finally, $d_{W_{p}}(\mu,\Phi^{-1}(\nu_{1}^{\prime}))=d_{W_{p}}(\mu,\mu^{\prime})=d_{W_{p}}(\mu,\Phi^{-1}(\nu_{2}^{\prime}))$ contradicts the fact that $\mu^{\prime}$ is the unique minimizer. This contradiction guarantees that $\Phi(\mu)\in\mathcal{F}$.
Figure 4. Illustration of the final step leading to a contradiction. Dashed lines represent equal distances.
After this brief sketch of the proof, we turn to the details.
If $\Phi(\mu)\not\in\mathcal{F}$, then there exist two points $z,z^{\prime}\in\mathrm{supp}(\Phi(\mu))$ such that their projections onto either $L_{+}$ or $L_{-}$ coincide. Indeed, if there are no such points, then all points of the support of $\Phi(\mu)$ project to different points of $L_{+}$ and $L_{-}$. Since
${P_{L_{-}}}_{\#}(\Phi(\mu))=\sum_{i=1}^{N}a_{i}\delta_{P_{L_{-}}(x_{i})},\ \ {P_{L_{+}}}_{\#}(\Phi(\mu))=\sum_{i=1}^{N}a_{i}\delta_{P_{L_{+}}(x_{i})}$
and $a_{i}\neq a_{j}$ ($1\leq i\neq j\leq N$), this implies that $\Phi(\mu)=\sum_{i=1}^{N}a_{i}\delta_{x_{i}}=\mu$, which leads to a contradiction. Without loss of generality, we can assume that this common projection is $P_{L_{+}}(x_{1})\in L_{+}$, i.e., $P_{L_{+}}(z)=P_{L_{+}}(z^{\prime})=P_{L_{+}}(x_{1})$, and for some $1\leq j_{1}\neq j_{2}\leq N$ we have $z=z_{1,j_{1}},z^{\prime}=z_{1,j_{2}}$ and $a_{1,j_{1}}>0,a_{1,j_{2}}>0$. Using this observation, we construct the measures $\mu^{\prime},\nu_{1}^{\prime},\nu_{2}^{\prime}$ as follows. We take a point $x^{\prime}\in L_{+}$ such that $c_{0}:=d_{m}(P_{L_{+}}(x_{1}),x^{\prime})<c/2$. Let us denote the elements of
$\Big{(}(P_{L_{+}})^{-1}(\{P_{L_{+}}(x^{\prime})\})\Big{)}\cap\Big{(}(P_{L_{-}})^{-1}(\{P_{L_{-}}(x_{1}),\ldots,P_{L_{-}}(x_{N})\})\Big{)}$
by $z_{0,j}$ ($1\leq j\leq N$) so that $P_{L_{+}}(z_{0,j})=x^{\prime}$ and $P_{L_{-}}(z_{0,j})=P_{L_{-}}(x_{j})$. We will also use the notation $x_{0}=z_{0,1}$. For every $0\leq i^{\prime}\leq N,1\leq j,j^{\prime}\leq N$
$d_{m}(z_{0,j},z_{1,j})<d_{m}(z_{i^{\prime},j^{\prime}},z_{1,j})$
if $(i^{\prime},j^{\prime})\notin\{(0,j),(1,j)\}$. To see this, observe that by construction, $d_{m}(z_{0,j},z_{1,j})=c_{0}$. If $i^{\prime}\neq 0$ and $(i^{\prime},j^{\prime})\neq(1,j)$, then we have $d_{m}(z_{i^{\prime},j^{\prime}},z_{1,j})>c$ by definition. If $i^{\prime}=0$ and $j^{\prime}\neq j$, then using the reverse triangle inequality, we have $d_{m}(z_{0,j^{\prime}},z_{1,j})\geq d_{m}(z_{1,j^{\prime}},z_{1,j})-d_{m}(z_{0,j^{\prime}},z_{1,j^{\prime}})\geq c-c_{0}>c_{0}.$ Let us fix a weight $a$ satisfying $0<a<\min\{a_{1,j_{1}},a_{1,j_{2}}\}<a_{1}$. Now, we consider the following measures
$\mu^{\prime}=a\delta_{x_{0}}+(a_{1}-a)\delta_{x_{1}}+\sum_{i=2}^{N}a_{i}\delta_{x_{i}},$
$\nu_{1}^{\prime}=a\delta_{z_{0,j_{1}}}+(a_{1,j_{1}}-a)\delta_{z_{1,j_{1}}}+a_{1,j_{2}}\delta_{z_{1,j_{2}}}+\sum_{j=1,\\ j\neq j_{1},j_{2}}^{N}a_{1,j}\delta_{z_{1,j}}+\sum_{i=2}^{N}\sum_{j=1}^{N}a_{i,j}\delta_{z_{i,j}},$
$\nu_{2}^{\prime}=a\delta_{z_{0,j_{2}}}+a_{1,j_{1}}\delta_{z_{1,j_{1}}}+(a_{1,j_{2}}-a)\delta_{z_{1,j_{2}}}+\sum_{j=1,\\ j\neq j_{1},j_{2}}^{N}a_{1,j}\delta_{z_{1,j}}+\sum_{i=2}^{N}\sum_{j=1}^{N}a_{i,j}\delta_{z_{i,j}}.$
Obviously, $\mu^{\prime},\nu_{1}^{\prime}$, and $\nu_{2}^{\prime}$ are probability measures satisfying $\mathcal{R}(\mu^{\prime})=\mathcal{R}(\nu_{1}^{\prime})=\mathcal{R}(\nu_{2}^{\prime})=:\mathcal{R^{\prime}}$, namely
$\mathcal{R^{\prime}}=\left(a\delta_{x^{\prime}}+(a_{1}-a)\delta_{P_{L_{+}}(x_{1})}+\sum_{i=2}^{N}a_{i}\delta_{P_{L_{+}}(x_{i})},\sum_{i=1}^{N}a_{i}\delta_{P_{L_{-}}(x_{i})}\right).$
Our next step is to prove that:
(5) $d_{W_{p}}(\mu,\mu^{\prime})=d_{W_{p}}(\Phi(\mu),\nu^{\prime}_{1})=d_{W_{p}}(\Phi(\mu),\nu^{\prime}_{2})=a^{\frac{1}{p}}c_{0},$
moreover, $\mu^{\prime}$ satisfies the following uniqueness property:
(6) $\text{if}\ d_{W_{p}}(\mu,\xi)=a^{\frac{1}{p}}c_{0},\ \text{and}\ \mathcal{R}(\xi)=\mathcal{R}^{\prime}\ \text{then}\ \xi=\mu^{\prime}.$
Equations (5) and (6) together will justify relations (3) and (4).
In order to show (5) note that if $\xi_{1},\xi_{2}$ are finitely supported probability measures with supports in a discrete set $P$, then
$d_{W_{p}}(\xi_{1},\xi_{2})=\min_{\Pi\in C(\xi_{1},\xi_{2})}\big{(}\sum_{(u,v)\in P\times P}d_{m}^{p}(u,v)\cdot\Pi(u,v)\big{)}^{\frac{1}{p}}.$
The proof of each of the equations in (5) is similar, therefore we will only prove the equality $d_{W_{p}}(\mu,\mu^{\prime})=a^{\frac{1}{p}}c_{0}$. Notice first, that since every transport plan must move a total weight of at least $a$ to $x_{1}$ from the support points of $\mu^{\prime}$, we have that
$d_{W_{p}}(\mu,\mu^{\prime})\geq\big{(}\sum_{i=0}^{N}d_{m}^{p}(x_{i},x_{1})\Pi(x_{i},x_{1})\big{)}^{\frac{1}{p}}\geq(d_{m}^{p}(x_{0},x_{1})a)^{\frac{1}{p}}=a^{\frac{1}{p}}c_{0},$
since $d_{m}^{p}(x_{i},x_{1})\geq d_{m}^{p}(x_{0},x_{1})$ and $\sum_{i=0}^{N}\Pi(x_{i},x_{1})\geq a$. On the other hand, if we move weight $a$ directly from $x_{0}$ to $x_{1}$, we exactly get that the cost of this transport plan is $a^{\frac{1}{p}}c_{0}$. Now we turn to the proof of (6). Let us suppose that we have a probability measure $\xi$ with $\mathcal{R}(\xi)=\mathcal{R}^{\prime}$. Then $\xi$ can be written in the form
$\xi=\sum_{i=0}^{N}\sum_{j=1}^{N}b_{i,j}\delta_{z_{i,j}},$
such that
$\sum_{j=1}^{N}b_{0,j}=a,\ \ \sum_{j=1}^{N}b_{1,j}=a_{1}-a,\ \ \sum_{j=1}^{N}b_{i,j}=a_{i}\ (2\leq i\leq N),\ \ \sum_{i=0}^{N}b_{i,j}=a_{j}.$
Again, every transport plan must move a total weight of at least $a$ to $x_{1}$ from the support points of $\xi$. Hence, we get that again
(7) $d_{W_{p}}(\xi,\mu)\geq\big{(}\sum_{i=0}^{N}\sum_{j=1}^{N}d_{m}^{p}(z_{i,j},x_{1})\Pi(z_{i,j},x_{1})\big{)}^{\frac{1}{p}}\geq(d_{m}^{p}(x_{0},x_{1})a)^{\frac{1}{p}}=a^{\frac{1}{p}}c_{0}.$
Let us recall that $d_{m}(x_{0},x_{1})<d_{m}(z_{i,j},x_{1})$, if $(i,j)\notin\{(0,1),(1,1)\}$. (Note that $z_{0,1}=x_{0}$, $z_{1,1}=x_{1}$.) Therefore, equality holds in (7) if and only if all transport occurs between $x_{0}$ and $x_{1}$ with weight $a$. This implies that $\xi-a\delta_{x_{0}}=\mu-a\delta_{x_{1}}$ and hence $\xi=\mu^{\prime}$.
In the last step, we show that the existence of $\nu_{1}^{\prime}\neq\nu_{2}^{\prime}$ implies that $\mu^{\prime}$ is not a unique minimizer in relation (3). Indeed, since $\Phi^{-1}$ is an isometry preserving measures supported on $L_{+}$ and $L_{-}$ we have by Lemma 6 that $\mathcal{R}(\Phi^{-1}(\nu_{1}^{\prime}))=\mathcal{R}(\Phi^{-1}(\nu_{2}^{\prime}))=\mathcal{R}^{\prime}$. Furthermore, according to (5), we have
$d_{W_{p}}(\mu,\Phi^{-1}(\nu^{\prime}_{1}))=d_{W_{p}}(\Phi(\mu),\nu^{\prime}_{1})=a^{\frac{1}{p}}c_{0}=d_{W_{p}}(\mu,\mu^{\prime}),$
and similarly,
$d_{W_{p}}(\mu,\Phi^{-1}(\nu^{\prime}_{2}))=d_{W_{p}}(\Phi(\mu),\nu^{\prime}_{2})=a^{\frac{1}{p}}c_{0}=d_{W_{p}}(\mu,\mu^{\prime}).$
Since $\Phi^{-1}(\nu_{1}^{\prime})\neq\Phi^{-1}(\nu_{2}^{\prime})$, this is a contradiction. ∎
@@ 3.1. Diagonal rigidity of $\mathcal{W}_{1}(\mathbb{R}^{2},d_{\mathrm{{m}}})$
    In words, this item means that $\mu$ admits a symmetrical with respect to every other measure.
@@
    In words, this item means that there is a measure $\eta$ aligned with both $(\mu_{1},\nu)$ and $(\mu_{2},\nu)$.
@@ 3.2. Diagonal rigidity of $\mathcal{W}_{p}(\mathbb{R}^{2},d_{\mathrm{{m}}})$ for $p>1$
In words, item (ii’) means that $\mu$ admits a symmetrical with respect to every other measure.
@@
To handle the case $\Phi_{t}$ for $t>0$, choose $\mu=\mu(0,1,0)=\frac{1}{2}\delta_{(-1,-1)}+\frac{1}{2}\delta_{(1,1)}$. Then $\Phi(\mu(0,1,0))=\mu(0,1,t)$. Now fix the Dirac measure $\delta_{(-1,0)}$ and calculate both $d_{W_{2}}(\delta_{(-1,0)},\mu(0,1,0))$ and $d_{W_{2}}(\delta_{(-1,0)},\mu(0,1,t))$. Again, if $\Phi$ would act like $\Phi_{t}$ on $L_{+}$, we should get the same result, since $\Phi$ is an isometry. The calculation shows that $d_{W_{2}}\big{(}\delta_{(-1,0)},\mu(0,1,0)\big{)}=\sqrt{\frac{5}{2}}$ and
@@ 3.3. Diagonal rigidity of $\mathcal{W}_{1}(Q,d_{\mathrm{{m}}})$
In this subsection, we consider the case when $X=Q=[-1,1]^{2}$ and $p=1$. Diagonal rigidity is achieved as a result of the following statements about measures that are supported on the sides, at the corners, and finally on the diagonals of $Q$. The first statement concerns measures supported on the opposite sides of the boundary of $Q$ and it is valid for all $p\geq 1$.
@@
Let $p\geq 1$ and $\Phi:\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})\to\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})$ be an isometry. If $\mu,\nu\in\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})$ are two probability measures whose supports lie on opposite sides of the closed unit ball $Q=[-1,1]^{2}$, then their isometric images $\Phi(\mu)$ and $\Phi(\nu)$ have the same property.
@@
with equality if and only if the supports $\mathrm{supp}(\mu)$ and $\mathrm{supp}(\nu)$ are contained in two opposite sides of $Q$. To see this, note that inequality (32) follows immediately from the definition of the Wasserstein metric $d_{W_{p}}$ and the fact that $d_{\mathrm{{m}}}(x,y)\leq 2$ for all $x,y\in Q$. Furthermore, if $\mathrm{supp}(\mu)$ and $\mathrm{supp}(\nu)$ are contained in two opposite sides of $Q$, then $d_{\mathrm{{m}}}(x,y)=2$ for all $x\in\mathrm{supp}(\mu)$ and $y\in\mathrm{supp}(\nu)$.
Let $\pi_{0}$ be an optimal coupling of $\mu$ and $\nu$. Since $\mathrm{supp}(\pi_{0})\subseteq\mathrm{supp}(\mu)\times\mathrm{supp}(\nu)$ we have that $d_{\mathrm{{m}}}(x,y)=2$ for any $(x,y)\in\mathrm{supp}(\pi_{0})$ and therefore
$d_{W_{p}}^{p}(\mu,\nu)=\int\limits_{Q\times Q}d_{\mathrm{{m}}}^{p}(x,y)~{}\mathrm{d}\pi_{0}(x,y)=2^{p}.$
To show the converse, let us assume that $d_{W_{p}}^{p}(\mu,\nu)=2^{p}$. Then for all couplings $\pi\in C(\mu,\nu)$ we have
$2^{p}=d_{W_{p}}^{p}(\mu,\nu)=\int\limits_{Q\times Q}d_{\mathrm{{m}}}^{p}(x,y)~{}\mathrm{d}\pi(x,y)\leq\int\limits_{Q\times Q}2^{p}~{}\mathrm{d}\pi(x,y)=2^{p},$
thus $d_{\mathrm{{m}}}(x,y)=2$ for $\pi$ almost all $(x,y)$. This applies for $\pi=\mu\otimes\nu$, and so $\mu$ and $\nu$ must be concentrated on the opposite sides of $Q$.
The statement of the lemma is now an immediate consequence of this claim. Indeed, assume that $\mathrm{supp}(\mu)$ and $\mathrm{supp}(\nu)$ are contained in two opposite sides of $Q$. Then we have $d_{W_{p}}^{p}(\mu,\nu)=2^{p}$. Since $\Phi:\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})\to\mathcal{W}_{p}(Q,d_{\mathrm{{m}}})$ is an isometry, we have $d_{W_{p}}^{p}(\Phi(\mu),\Phi(\nu))=2^{p}.$ But, then the supports of the two measures $\Phi(\mu)$ and $\Phi(\nu)$ must be contained in two opposite sides of $Q$. ∎
@@
$V=\{\delta_{x_{1}}=\delta_{(-1,-1)},\delta_{x_{2}}=\delta_{(1,-1)},\delta_{x_{3}}=\delta_{(1,1)},\delta_{x_{4}}=\delta_{(-1,1)}\}.$
@@
It is clear by Lemma 13 that all other measures $\Phi(\delta_{x_{i}})$ for $i=2,3,4$ must be supported on the opposite side of $S_{1}$, that is $S_{3}.$ But the mutual distance of any pair of these measures must be equal to $2$ which shows that any two of these three measures must be lying on opposite sides again, which is $S_{2}$ and $S_{4}$. But there are only two possibilities for measures with support in $S_{2}\cap S_{3}$ and $S_{4}\cap S_{3}$, namely the two Dirac masses on the vertices of $S_{3}$ which gives a contradiction.